A flywheel with a diameter of 1.0 m and initially rotating at 500 rpm, rotate at 1000 rpm after 20 s. Assuming a constant angular acceleration, find the value of the angular acceleration.





1
Expert's answer
2021-02-10T08:35:07-0500

"500\ rpm=8.33\ s^{-1}"


"1000\ rpm=16.67\ s^{-1}"


"\omega=\omega_0+\epsilon t \to \epsilon=\frac{\omega-\omega_0}{t}=\frac{2\pi\nu-2\pi\nu_0}{t}="


"=\frac{2\pi\cdot 16.67-2\pi\cdot8.33}{20}=2.62\ (rad/s^2)" . Answer